HUBBLE SPACE TELESCOPE DAILY REPORT #2210 PERIOD COVERED: 0000Z (UTC) 09/16/98 - 0000Z (UTC) 09/17/98 Daily Status Report as of 260/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7984 (HDF Limiting Magnitude/Resolution Extension and Object Variability Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations to be used for HDF limiting magnitude/resolution extension and object variability search. This will be a search for supernovae at high Z in the HDF. A bias and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Five Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Eight Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 7628 (WF/PC-2 Cycle 7 Photometric Characterization) The WF/PC-2 was used to make observations of globular cluster NGC-2100 as part of a photometric characterization calibration observation. This is to determine whether any changes in the zeropoint, the spatial dependence of the zeropoint, or contamination rates have occurred. The observation was executed as scheduled, and no anomalies were reported. 1.8 Completed Two Sets of STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.9 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star CD-40D5404. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/1/2/3 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around the nearby, young main sequence star HD21703. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.12 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-1553 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.13 Completed Two Sets of NIC/2 7835 (A Search for Superplanets Embedded in Beta Pic & Vega-like Circumstellar Disks) The Near Infrared Camera (Camera Two) was used to make observations of the circumstellar disk around star HD39060 in order to search for superplanets. The discovery of a circumstellar disk orbiting Beta Pictoris as well as the discovery of other similar IR-excess Vega-like systems have provided compelling evidence that our solar system, or at least many of the processes which are thought to have created it, may be far from unique. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.15 Completed NIC/1/2/3 7830 (IR Spectra for Known-Mass M Dwarfs) The Near Infrared Camera (Camera Three) was used to make infrared spectrographic observations of M-Class dwarf star GJ2005ABC. Over the last decade there has been considerable debate on two major issues in the low-mass star literature: {1} the effective temperature scale for M dwarfs and its relation to mass and luminosity and {2} the conversion from an absolute magnitude in some passband to a bolometric magnitude. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Two Sets of NIC/2 7441 (A Search for Zodiacal Dust around Bright Nearby Stars) The Near Infrared Camera (Camera Two) was used to make observations of the zodiacal lights around the bright nearby star Tau-Cet. Dust particles around a main sequence star are quickly removed by Poynting-Robertson drag, therefore, the existence of such dust is an almost certain signature of a current presence of a population of larger bodies in orbit around the star. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 6619 (High Resolution Images of QSO Lyman Alpha Absorbing Galaxies) The WF/PC-2 was used to make several images of galaxy field (Field 0407) in order to obtain high resolution observations of the Lyman Alpha absorbing galaxies within this field-of-view. The purpose is to directly identify the objects responsible for low and intermediate redshift QSO Lyman-Alpha systems. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed WF/PC-2 6681 (Evolution of Pre-Main Sequence Circumstellar Nebulosity) The WF/PC-2 was used to make observations of V819-Tau to study the evolution of pre-main sequence circumstellar nebulosity. Initial HST results from imaging of disks around pre-main sequence stars provide compelling justification to observe a broader sample of young stars. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 14 Successful : 14 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 30 Successful : 30 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. As described in HSTAR 6676, the STIS MCE-2 reset at 259/131644Z, outside an SAA interval and while the high voltage was off. ROP NS-12 was executed at 259/134314Z to clear the STIS error counter. Recovery will be accomplished via SMS commanding at 264/022938Z, the time of the next MAMA-2 high voltage on. There were two instances of NICMOS EDAC single bit errors, at 259/145556Z and at 259/224056Z. In each case, NS-11 was executed to dump memory and to reset the NICMOS error counter. After the final occurrence, the parameter limit was changed to 16. HSTARs 6677 and 6678 were written. The HST IP was again the prime system from approximately 259/1200Z until approximately 259/2000Z with no reported problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH